OpenStudy (anonymous):

What is \[7^0 + \sqrt{7^1 + \sqrt{7^2 + \sqrt{7^4 + \sqrt{7^8 + \sqrt{7^{16} + \sqrt{7^{32} + \sqrt{7^{64} + ...}}}}}}}\] equal to?

OpenStudy (jamesj):

Might as well generalize the problem now to any positive integer n

OpenStudy (anonymous):

\[y = 7^0 + \sqrt{7^1 + \sqrt{7^2 + \sqrt{7^4 + \sqrt{7^8 + \sqrt{7^{16} + \sqrt{7^{32} + \sqrt{7^{64} + ...}}}}}}}\] \[y = 1 + \sqrt{y}\] maybe I am not sure

OpenStudy (jamesj):

And I notice agdgd that you like to ask questions but I haven't seen you solve many yet. So let's see something. Show that \[1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + ...}}}}}}}\] converges.

OpenStudy (jamesj):

@ishaan: see agdgd's earlier post.

OpenStudy (anonymous):

\[n^0 + \sqrt{n^1 + \sqrt{n^2 + \sqrt{n^4+\sqrt{n^8+\sqrt{n^{16}+...}}}}}=1 + \sqrt{n}*\phi = \frac{\sqrt{5n} + \sqrt{n}+2}{2}\]

OpenStudy (anonymous):

Oh Okay thanks

OpenStudy (anonymous):

I'm not too good at solving questions :(

OpenStudy (anonymous):

so we use induction?

OpenStudy (anonymous):

oh well somebody else solved this problem already: http://www.math.washington.edu/~morrow/336_11/papers/brian.pdf

OpenStudy (anonymous):

better than I ever could

OpenStudy (anonymous):

so we do use induction :-D

OpenStudy (anonymous):

hmm is it \[y = 1 + \sqrt{7y}\]

OpenStudy (anonymous):

oh no sorry mistake my bad

OpenStudy (anonymous):

so we define \[\sqrt{a+\sqrt{a+\sqrt{a+...}}}\]as \[t_n(a)\]where\[t_1(a) = \sqrt{a}\] and \[t_{n+1}(a)=\sqrt{a +t_n}\]

OpenStudy (anonymous):

how does induction work here?

OpenStudy (anonymous):

step 1: show that t(1) is convergent that's easy enough

OpenStudy (jamesj):

The paper you link to glosses the proof that it converges, just saying it's "easy".

OpenStudy (anonymous):

I hope it is easy enough for me :(

OpenStudy (anonymous):

so I think now I have to show that \[t_{n+1}(1)=\sqrt{1 +t_n}\] is convergent?

OpenStudy (anonymous):

so we can say that as n approaches infinity, t_n+1 = t_n ?

OpenStudy (anonymous):

\[t_n(1)^2 - t_n(1) -1 = 0\]

OpenStudy (jamesj):

You show it's convergent by showing that the sequence (tn) is monotonically increasing and bounded above.

OpenStudy (anonymous):

so I find the derivative?

OpenStudy (jamesj):

Nope.

OpenStudy (jamesj):

Show that \( t_{n+1} > t_n \)

OpenStudy (anonymous):

using induction?

OpenStudy (jamesj):

and show there exists a number B such that \( t_n < B \) for all \( n \).

OpenStudy (anonymous):

Differentiate it

OpenStudy (jamesj):

nope. It's not a continuous function.

OpenStudy (anonymous):

Oh sorry

OpenStudy (anonymous):

so t_0(1) = 1, t_1(1) = \sqrt(2), t_2(1) = \sqrt{1 + \sqrt{2}} then what?

OpenStudy (jamesj):

1. First show that \( \varphi\) is an upper bound of the sequence by showing that \[ x < \varphi \ \implies \sqrt{1+x} < \varphi \ \ \ --- (*)\] If you prove that, then as we know that \( 1 < \varphi \), it will follow by the PMI that \( t_n(1) < \varphi \) for all \( n \geq 1 \). 2. Then show that \( (t_n) \) is monotonically increasing. It will turn out the result of the 1. will be critical. It is then time to use one of the most basic and important theorems in analysis: If \( (x_n) \) is a monotonically increasing sequence of real numbers bounded above, then \( (x_n) \) converges. Don't try and prove this theorem; just use it. You can find the proof in Spivak or any other first book on analysis.

OpenStudy (anonymous):

I think step 1 can be done by showing something about phi.

OpenStudy (jamesj):

Try it for once instead of second guessing: Suppose \[ x < \varphi \] then \[ \sqrt{1+x} < \sqrt{1 + \varphi} = ... \]

OpenStudy (anonymous):

it follows from \[\phi = \sqrt{\phi + 1}\]that \[\sqrt{1+x}<\phi\]

OpenStudy (anonymous):

but then what do I do next?

OpenStudy (anonymous):

so I simply plug in t_n?

OpenStudy (jamesj):

Prove that \[ \varphi = \sqrt{1 + \varphi} \]

OpenStudy (jamesj):

You show it.

OpenStudy (anonymous):

how could I do that without knowing what phi is?

OpenStudy (jamesj):

We're starting this chain of logic with \[ \varphi = \frac{\sqrt{5}+1}{2} \]

OpenStudy (anonymous):

I think thinking about golden ratios and all for this problem is not going to help, golden ratio like pi keeps popping in various things in mathematics like this one ..

OpenStudy (anonymous):

:(

OpenStudy (jamesj):

Actually, it's pretty much essential for showing that this sequence \( (t_n) \) converges.

OpenStudy (anonymous):

I think agd problem is he/he is somehow trying to learn everything at once may be that's causing cognitive overflow ?

OpenStudy (anonymous):

Maybe i simply don't get induction

OpenStudy (anonymous):

or maybe i'm not putting in the effort

OpenStudy (anonymous):

i'm too used to uncle Google and the Spivak answer book

OpenStudy (jamesj):

Let \[ \varphi = \frac{\sqrt{5}+1}{2} \] and define a sequence \( (t_n) \) by \[ t_1 = 1, \ \ t_{n+1} = \sqrt{1 + t_n} \ \ \hbox{ for } n \geq 1\] Let P(n) be the statement \( t_n < \varphi \). P(1) is the statement \( 1 < \varphi \). This is clearly true as \( 2 < \sqrt{5} + 1 \) because \( 1 < \sqrt{5} \) because \( 1 < 5 \). Now we want to show that for \( k \geq 1 \), the statement \( P(k) \) implies \( P(k+1) \). Suppose P(k). I.e., \( t_k < \varphi \). Then \( \ \ \ \ \ t_{n+1} \) \[ = \sqrt{1 + t_n} \] \[ < \sqrt{1 + \varphi} \ \ \ \hbox{by hypothesis, i.e., by the statement P(k)} \] Now \[ \varphi = \sqrt{1 + \varphi} \] iff \[ \varphi^2 = 1 + \varphi \] iff \[ \left( \frac{\sqrt{5}+1}{2} \right)^2 = 1 + \frac{\sqrt{5}+1}{2} \] iff \[ \frac{6 + 2\sqrt{5}}{4} = \frac{3 + \sqrt{5}}{2} \] which is true. Therefore P(k) => P(k+1) for an arbitrary \( k \geq 1 \). Hence, as P(1) is also true, it follows by the Principle of Mathematical Induction that P(n) is true for all \( n \geq 1 \).

OpenStudy (jamesj):

Now we have shown that \( (t_n) \) is bounded above. This was the statement P(n). Now YOU show that \( (t_n) \) is an strictly increasing sequence.

OpenStudy (anonymous):

Great answer James but may be you are expecting too much out of a guy who probably didn't studied real analysis or probably analysis before ? ;)

OpenStudy (anonymous):

We need to show that \[t_{n+1} - t_n > 0 \text{ for all n}\geq1\]squaring both sides gives us \[(1 + t_n) - 2(t_{n+1} * t_n) + t_n^2 > 0\]

OpenStudy (anonymous):

This will keep me up all day :(

OpenStudy (anonymous):

or how about \[1 + t_n > t_n^2\]iff\[-(t^2 -t_n -1) > 0\]

OpenStudy (anonymous):

t_n^2 * oops

OpenStudy (anonymous):

\[-(t_n^2 -t_n -1) > 0\]iff \[t_n <\frac{1+\sqrt{5}}{2}\]this is true since \[t_{n+1} <\phi\]iff\[\sqrt{1 + t_n}<\sqrt{1 + \phi}\]iff\[t_n < \phi\] so I think we've successfully shown that the sequence t_n is monotonically increasing

OpenStudy (jamesj):

why is this step true?

OpenStudy (anonymous):

if you find the roots of the polynomial, it turns out that \[\frac{1+\sqrt{5}}{2}\]is the upper bound

OpenStudy (jamesj):

ok. You're not getting an A in analysis any time soon, but you're not getting an F today either. :-)

OpenStudy (anonymous):

:( it took me an hour to figure that one out :-P

OpenStudy (jamesj):

btw, when did you change into a black man? or were you always a black man?

OpenStudy (anonymous):

I've always been george agdgdgwngo

OpenStudy (jamesj):

then how come when I first knew you you had a picture of a young girl as your profile pick?

OpenStudy (anonymous):

I've been a young girl once.

OpenStudy (jamesj):

by which I mean 19 years old or so.

OpenStudy (jamesj):

ok then.

OpenStudy (anonymous):

How do I proceed from all this?

OpenStudy (anonymous):

If I ever want to master the 'proof by induction'

OpenStudy (jamesj):

like I told you when you first asked about induction, you need to write a few out, get them corrected, like just about everything in mathematics. Failing that, read carefully written induction proofs, like the one I've written in this thread, or in good books like Spivak.

OpenStudy (jamesj):

If you want to practice, write out a proof for by induction for this identity: 1!1 + 2!2 + 3!3 + ... + n!n = (n+1)! - 1. Write out the proof in four sections, beginning exactly as follows: "Let \(P(n)\) be the statement ..." "First, the statement \( P(1) \) is true because ..." "Next we want to show that for \( k \geq 1 \ P(k)\implies P(k+1)\). \( P(k) \) says that ..." "Hence having shown that \(P(1)\) is true and that for \( k \geq 1 \), \( P(k)\implies P(k+1) \), it follows by the Principle of Mathematical Induction that ..."

OpenStudy (anonymous):

\[\small{\text{Let }}P(n)\small{\text{ be the following statement:}}\]\[1!1 + 2!2 + 3!3 + ... + n!n = (n+1)! - 1\]\[\small\text{Hence,}\] P(1)\[\text{ is the statement}\]\[1!1 = (1+1)!-1\]iff\[1=1\]which is true \[\text{Next we want to show that for }k \geq1, \ P(k)\implies P(k+1).\text{}\]\[\text{If all of }P(1), P(2)...P(k)\text{ are true, then in particular, }P(k)\text{ is true.}\]\[P(k)\text{ says that:}\]\[1!1 + 2!2 + 3!3 + ... + k!k = (k+1)! - 1\]adding \[(k+1)!*(k+1)\]to both sides gives us \[1!1 + 2!2 + 3!3 +...+ k!k + (k+1)!(k+1) = (k+2)!-1\]which shoes that P(k+1) is also true. Hence having shown that P(1) is true and that for k≥1, P(k)⟹P(k+1), it follows by the Principle of Mathematical Induction that P(n) is true for all n \geq 1

OpenStudy (anonymous):

Am I mistaken?

OpenStudy (jamesj):

Yes, four small and one large problems, most of which are the second step of the proof, "P(k) => P(k+1)" * In the first step, it would be better to say that 1!1 = 2! - 1 iff 1 = 2 - 1 * We are NOT assuming all of P(1), P(2), ... P(k-1), P(k). We are only using the hypothesis P(k). * In this step we do not say P(k) is true, as we don't know that yet. We say "suppose P(k)"; we are using that statement as a hypothesis. *** This is the most serious issue: In the heart of the proof, after you've added (k+1)!(k+1) to both sides, you skipped the critical calculations in the RHS. Why is it true that (k+1)! - 1 + (k+1)!(k+1) = (k+2)! - 1 ? I'm not asking you write it out for me now; I'm simply saying this is the most important part of the proof and you skipped over it. * At the end of this step, we do not say "hence P(k+1) is true", as we still don't know that yet. All we have shown is that P(k) => P(k+1). Remember that implication can be true without P(k) being true.